OpenStudy (anonymous):

simplify (2 exponent) 3

OpenStudy (anonymous):

2^3 = 2*2*2 = ?

OpenStudy (anonymous):

whats 2*2 =

OpenStudy (anonymous):

4

OpenStudy (anonymous):

now 4*2 =

OpenStudy (anonymous):

6

OpenStudy (anonymous):

4*2 is not 6

OpenStudy (anonymous):

whats 4+4 =

OpenStudy (anonymous):

lol....8

OpenStudy (anonymous):

so whats 4*2 =

OpenStudy (anonymous):

8

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

but the answers that i have to choose from dnt show 8

OpenStudy (anonymous):

2 to the exponent of 3 is 8 just write 8 as your answer

OpenStudy (anonymous):

it shows 44 55 64 and 74

OpenStudy (anonymous):

well 2 cubed is 8

OpenStudy (anonymous):

well i found out how to do another problem

OpenStudy (anonymous):

2^3=8

OpenStudy (anonymous):

the problem i have is(2 with a small 2 above it ) 3

OpenStudy (anonymous):

they are algebraic problems

OpenStudy (ajprincess):

Is ur question \((2^2)^3\)?

OpenStudy (anonymous):

the 2 in parenthesis is right and there is a normal size 3 also

OpenStudy (ajprincess):

Can u plz draw ur question?

OpenStudy (anonymous):